OpenStudy (anonymous):

Which of the following is equal to the fraction below? (5/9)^8 A. 8 *(5/9) B. 5/9^8 C. 5^8/9^8 D. 5^8/9

OpenStudy (bibby):

\(\large (\frac{a}{b})^c=(\frac{a^c}{b^c})\)

OpenStudy (anonymous):

so it would be C then? 5^8/9^8?

OpenStudy (bibby):

yeah, probably, maybe

OpenStudy (thelittlemerpmaid):

C sounds right, but if it's not C, then it's B. That was my first guess.

OpenStudy (bibby):

if you don't believe me, wolfram confirms http://www.wolframalpha.com/input/?i=5%5E8%2F9%5E8%3D%285%2F9%29%5E8

OpenStudy (anonymous):

I do!

OpenStudy (bibby):

it's not that bad to not believe what I say question everything. later dudes